OpenStudy (anonymous):

Solve: 2 (Radical x) + 8 = -2

OpenStudy (allank):

I assume the (Radical x) is \[\sqrt{x}\] ?

OpenStudy (anonymous):

\[2 \sqrt{x} = -10\] Squaring both the sides, \[4x = 100\] Divide by 4 both the sides and tell me what you got..

OpenStudy (anonymous):

don't you have to divide by 2 first before you square both sides?

OpenStudy (anonymous):

You can do that also.. Answer will remain the same.. According to you: Divide by 2 both the sides: \[\large \sqrt{x} = -5\] Squaring both the sides: \[\large \color{green}{x = 25}\]

OpenStudy (anonymous):

And According to me: \[\large 4x = 100\] Divide by 4 both the sides: \[\large x = \frac{\cancel{100}^{25}}{\cancel{4}} = 25\]

OpenStudy (anonymous):

okay thank you so much for the help!

OpenStudy (anonymous):

Welcome dear..